find an equation for the line with the given properties. Express the equation in slope-intercept form. Slope=4 containing points (-7,1)
Answer by jim_thompson5910(35256)


y = mx+b

1 = 4(-7)+b

1 = -28+b

1+28 = b

29 = b

b = 29

So the equation of the line is y = 4x+29


Let me know if you need more help or if you need me to explain a step in more detail.
Feel free to email me at jim_thompson5910@hotmail.com
or you can visit my website here: http://www.freewebs.com/jimthompson5910/home.html

Thanks,

Jim